# Comparison Principles for Stochastic Volterra equations

### v1

@@ Abstract.
In this work, we establish a comparison principle for stochastic Volterra equations with respect to the initial condition and the drift $b$ applicable to a wide class of Volterra kernels and driving forces $g$ that may be singular in zero. For completely monotone Volterra kernels such a result holds without any further restrictions, while for not completely monotone kernels it is shown that such a principle fails unless the drift is additionally monotone. As a side-product of our results, we also complement the literature on the weak existence of continuous nonnegative solutions. This covers the rough Cox-Ingersoll-Ross process with singular initial conditions.
Key words and phrases:
stochastic Volterra process; comparison principle; rough Cox-Ingersoll-Ross process; splitting method
@@ 1. Introduction
In many applications, one assumes that the Volterra kernel is given by the fractional Riemann-Liouville kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha\in(1/2,2)$. Such equations have gained increased attention in mathematical Finance for the modelling of rough volatility, see [8, 14, 15, 16, 17, 23]. The weak existence of a solution was established in [4] whenever $b,\sigma$ are continuous with linear growth, $g$ constant and $K$ satisfies an additional regularity condition given below. An extension to processes with jumps and more general $g$ is discussed in [1]. When $b$ and $\sigma$ are Lipschitz continuous, the existence of a unique strong solution was shown in [2, Theorem A.1], see also [9, 31]. For non-Lipschitz coefficients, the pathwise uniqueness of solutions is more subtle and yet not fully understood. For the fractional Riemann-Liouville kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha\in(1/2,1)$, the method proposed in [30] gives the existence and uniqueness of strong solutions whenever $\sigma$ is $\eta$-Hölder continuous and $\alpha\eta>\frac{1}{2}$. An extension of this method towards general drifts and non-convolution equations was recently discussed in [32]. Most recently, in [26], the author proves the weak existence and uniqueness in law for (1) provided that $b,\sigma$ are uniformly continuous, $\sigma$ is non-degenerate, and some technical condition on the modulus of continuity holds. In contrast to the existing literature, see also [32, 33], in this work we focus on continuous solutions where $g$ is not regular but belongs to $g\in L_{\mathrm{loc}}^{p}(\mathbb{R}_{+})\cap C((0,\infty))$ with an additional growth condition at the origin. Since such a case is yet not covered by the literature, we first provide in Theorem 2.6 a general result on the weak existence of continuous solutions for continuous coefficients with linear growth. In the main part of this work, we study comparison principles for solutions of (1). More precisely, given initial data $(g_{i},K,b_{i},\sigma)$ with $i=1,2$, we provide sufficient conditions such that the corresponding solutions $X^{1},X^{2}$ satisfy
@@
While such comparison principles are well-known for classical ordinary differential equations (see [38]) and for Markovian stochastic equations (see [27] for the continuous case and [21] for the case of continuous-state branching processes) much less is known for their Volterra counterparts. For instance, comparison principles for deterministic Volterra equations have been studied in [25, Chapter 12.1], while the first results on stochastic Volterra equations were obtained in [18, 36] for general drifts and restrictive assumptions on the diffusion which exclude our convolution setting (1). Based on the splitting method, Alfonsi was studied in [6] a numerical approximation of stochastic Volterra equations applicable to $g(t)=x$ constant, globally Lipschitz continuous coefficients $b,\sigma$, and $K$ being either non-negativity preserving with $K(0)>0$ or completely monotone. The latter work includes, up to our knowledge, the first general comparison principle for equations of the form (1). This splitting method has been also recently applied in [7] to prove the existence and uniqueness of nonnegative solutions of stochastic Volterra equations with jumps for non-Lipschitz continuous coefficients and regular Volterra kernels $K\in C^{2}(\mathbb{R}_{+})$. The reader interested in other numerical schemes for stochastic Volterra processes may consult [3, 5] and the references therein. In this work, we first extend Alfonsi’s numerical approximation [6] to non-constant $g$ with a particular focus on possibly singular functions, and time-dependent coefficients. Based on this approximation, we obtain a comparison principle that covers such cases including not necessarily Lipschitz continuous. Afterwards, we focus on kernels $K$ that are not completely monotone hence going beyond the framework discussed in [6]. In this generality, it is easy to see that for non-constant $g$ the comparison principle may fail (see the fractional Ornstein-Uhlenbeck process discussed in Example 2.5) unless further conditions on the driving function $g$ are imposed. The latter motivates the following definition.
Definition 1.1.
@@
Let $g_{1},g_{2}\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})\cap C((0,\infty))$ be such that
$g_{i}(t)=\widetilde{g}_{i}(t)+\int_{0}^{t}K(t-s)h_{i}(s)\,\mathrm{d}s,\qquad i=1,2,$
where $\widetilde{g_{1}},\widetilde{g}_{2}\in C((0,\infty))$ and $h_{1},h_{2}\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})$. We call $(g_{1},g_{2})$ admissible, if they are of the above form, $\widetilde{g}_{2}-\widetilde{g}_{1}\geq 0$ is non-decreasing on $(0,\infty)$, and $h_{2}-h_{1}\geq 0$.
This class of admissible functions is sufficiently rich to cover the most frequently used examples from applications including fractional kernels as illustrated below.
@@
Given the fractional Riemann-Liouville kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha\in(1/2,1)$. Define $g_{i}(t)=x_{i}\frac{t^{\gamma-1}}{\Gamma(\gamma)}$ with $\gamma>\alpha$ and $x_{1}\leq x_{2}$. Then $g_{1},g_{2}$ are admissible with
  1. (i)
    If $\gamma\in(\alpha,1)$, then $\widetilde{g}_{i}(t)=0$ and $h_{i}(t)=x_{i}\frac{t^{(\gamma-\alpha)-1}}{\Gamma(\gamma-\alpha)}$.
  2. (ii)
    If $\gamma\geq 1$, then $\widetilde{g}_{i}(t)=x_{i}\frac{t^{\gamma-1}}{\Gamma(\gamma)}$ and $h_{i}(t)=0$.
To illustrate our results, we consider the particular case of the fractional Ornstein-Uhlenbeck process given by
(2) $X_{t}=\frac{t^{\gamma-1}}{\Gamma(\gamma)}x+\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}\left(b+\beta X_{s}\right)\,\mathrm{d}s+\sigma\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}\,\mathrm{d}B_{s}.$
where $x,b,\beta\in\mathbb{R}$, $\sigma,\gamma>0$, and $\alpha\in(1/2,1)$. The precise formulation is given in Example 2.5. Furthermore, based on our weak existence result obtained in Theorem 2.6 combined with the comparison principle, in Theorem 2.8 we obtain the weak existence of nonnegative weak solutions of (1). In particular, in Example 2.9, we construct a weak nonnegative solution for the rough Cox-Ingersoll-Ross process
$X_{t}=g(t)+\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}\left(b+\beta X_{s}\right)\mathrm{d}s+\sigma\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}\sqrt{X_{s}}\,\mathrm{d}B_{s},\qquad t>0.$
for a large class of functions $g$ that may be singular in $t=0$. The latter extends [2] where $g$ was assumed to be regular in zero, and provides a simpler condition on the nonnegativity of the solutions that does not require bounds on the increments of the resolvent of the first kind. The first key obstacle when dealing with stochastic Volterra equations stems from the absence of the semimartingale property. Indeed, while the classical methods described in [27, Chapter VI] are deeply based on the Itô formula, in our setting an analogous approach is only applicable when $K$ is sufficiently regular in the sense that $K\in W^{1,2}_{\mathrm{loc}}(\mathbb{R}_{+})$. The latter excludes, e.g., the Riemann-Liouville kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha\in(1/2,3/2)$. To address this issue, we use the splitting method proposed in [6] to approximate the solution via solutions of classical stochastic differential equations. Observing that the approximation satisfies the comparison principle, we may then pass to the limit. This approach covers $K$ being either non-negativity preserving with $K(0)>0$ or completely monotone. In particular, it covers the fractional kernel with $\alpha\in(1/2,1)$. On the other side, if $K$ is neither non-negativity preserving nor completely monotone, a comparison principle is, in general, not valid as demonstrated in Example 3.1. The latter is, e.g., the case for the fractional kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha>1$. Such a failure of the comparison principle is not surprising since solutions of (1) are typically not Markov processes and hence it is not clear if the past trajectory entering the equation via $b,\sigma$ violates the ordering of solutions. To exclude such a possibility we suppose that the drift is additionally monotone. Then, we first consider the case of regular kernels $K\in W_{\mathrm{loc}}^{1,2}(\mathbb{R}_{+})$ for which a semimartingale representation is available and a modification of the classical comparison principle for SDEs can be used to deduce the comparison principle. Afterwards, we extend these results to general $K\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})\cap C^{1}((0,\infty))$ with $K^{\prime}\geq 0$ or $K^{\prime}\leq 0$ by approximation. This work is organised as follows. In Section 2 we treat the case of completely monotone kernels, where we first prove a comparison principle for Lipschitz continuous $b,\sigma$. Then we prove the existence of continuous weak solutions for continuous $b,\sigma$ with at most linear growth, and finally derive the corresponding comparison principle in this case. As an application, we conclude Section 2 with a result on the weak existence of continuous nonnegative solutions of (1). Afterwards, in section 3 we study similar results for the case of not completely monotone kernels. Some technical results on the existence, uniqueness, and sample path regularity of solutions to stochastic Volterra equations are collected in Appendix A, while the convergence of the splitting method is studied in Appendix B.
2. Non-negativity preserving case
In this section, we study the case where the Volterra kernel preserves nonnegativity as recently introduced and studied in [6].
@@ 2. Non-negativity preserving case
    (3) $\sum_{\ell=1}^{k}x_{\ell}K(t_{k}-t_{\ell})\geq 0,\ \ \forall k\in\{1,\dots,N\}\Longrightarrow\sum_{k=1}^{N}\mathbbm{1}_{\{t_{k}\leq t\}}x_{k}K(t-t_{k})\geq 0,\quad\forall t\geq 0.$
Here we use the convention $\sum_{\varnothing}\coloneqq 0$. Such a condition allows us to show that the Volterra kernel preserves the order structure subject to the discretisation of the process via the splitting method. Let us assume that the input data $(g,K,b,\sigma)$ of (1) satisfies throughout the entire paper the following regularity condition.
Assumption 2.2.
There exist $C>0$, $\xi\in[0,1]$ such that
$|b(t,x)|+|\sigma(t,x)|\leq(1+|x|)^{\xi}$
@@
    $g\in L^{q}_{\mathrm{loc}}(\mathbb{R}_{+})$ for some $q\in(2,\infty)$, $\xi\in(0,1]$, there exists $\eta>0$ such that $K\in L^{2+\eta}_{\mathrm{loc}}(\mathbb{R}_{+})$ and there exist $\gamma\in(0,1/2]$ with and for each $T>0$ a constant $c(T,\eta)$ such that for $h\in[0,1]$
    Moreover, the value $q$ satisfies
    (4) $q>2\xi\frac{1+\frac{1}{\eta}}{\gamma+\frac{1}{2}\frac{\eta}{2+\eta}}.$
The above assumption allows us to assume without loss of generality that for any solution $X$ of (1), $X-g$ is continuous, c.f. Proposition A.2. Note that such a condition is satisfied by the fractional Riemann-Liouville kernel as demonstrated in the next example.
Example 2.3.
Let $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha\in(1/2,1]$. Then $2+\eta<\frac{1}{1-\alpha}$, $\gamma=\alpha-1+\frac{1}{2+\eta}$. Hence (4) takes the particular form $q>2\xi\frac{1+1/\eta}{\alpha-1/2}$. By letting $\eta\nearrow\frac{2\alpha-1}{1-\alpha}$, we see that this condition is satisfied whenever
$q>\frac{\xi\alpha}{\left(\alpha-\frac{1}{2}\right)^{2}}.$
The following is our first result for non-negativity preserving kernels.
@@
Suppose that $(g_{i},K,b_{i},\sigma)$ with $i=1,2$ satisfy assumption 2.2, that $K\colon(0,\infty)\longrightarrow\mathbb{R}_{+}$ is continuous and preserves non-negativity, and let $g_{1},g_{2}$ be admissible such that there exists $\delta>0$ with
(5) $\int_{0}^{T}\overline{g}_{i}(t)^{2+\delta}(1+h_{i}(t)^{2})\,\mathrm{d}t$ $<\infty,\qquad i=1,2$
where $\overline{g}_{i}(t)=\sup_{s\in[t,T]}|g_{i}(s)|$. Let $b_{1},b_{2},\sigma\colon\mathbb{R}_{+}\times\mathbb{R}\longrightarrow\mathbb{R}$ be measurable and there exists $C>0$ such that for a.a. $t\geq 0$ and $x,y\in\mathbb{R}$,
@@
Finally, assume that the ordering condition $b_{1}(t,x)\leq b_{2}(t,x)$ holds. Let $X^{1},X^{2}$ be the unique strong solutions of
(7) $X_{t}^{i}=g_{i}(t)+\int_{0}^{t}K(t-s)b_{i}(s,X^{i}_{s})\,\mathrm{d}s+\int_{0}^{t}K(t-s)\sigma(s,X^{i}_{s})\,\mathrm{d}B_{s}$
for $i=1,2$. Then $\mathbb{P}[X^{1}_{t}\leq X^{2}_{t},\,t>0]=1$.
@@
First note that equation (7) has a unique strong solutions $X^{1},X^{2}$ with sample paths in $L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$ due to Proposition A.1. Moreover, given Assumption 2.2, it follows from Proposition A.2 that $X^{1}-g_{1}$, $X^{2}-g_{2}$ have a modification with continuous sample paths. Hence $X^{1},X^{2}$ are continuous on $(0,\infty)$. Finally, define
@@
Hence $\widetilde{b}_{1}\leq\widetilde{b}_{2}$ and (6) still holds with the same $C$, while the linear growth constant becomes $|\widetilde{b}_{1}(t,x)|+|\widetilde{b}_{2}(t,x)|\leq\widetilde{C}(t)(1+|x|)$ with $\widetilde{C}(t)=C+|h_{1}(t)|+|h_{2}(t)|$, so that $\widetilde{C}\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$. The proof is divided into three steps. Step 1: We introduce a similar splitting method to Alfonsi [6], but now applied for non-constant possibly singular at zero $g$ and time-dependent coefficients. Suppose first that $K$ is continuous with $K(0_{+})>0$. Let $T>0$, $N\in\mathbb{N}$ be fixed and $t_{k}=kT/N,$ $k\in\{0,1,\dots,N\},$ be an equidistant grid of $[0,T]$ with step size $T/N.$ Recursively, we define the process $(\widehat{X}^{i}_{t})_{t\in(0,T]}$ by setting
@@
Thus it suffices to prove that $\widehat{X}_{t}^{1}\leq\widehat{X}_{t}^{2}$ a.s. for $t\in(0,T]$. Step 2: We consider the case where $K$ is continuous and non-negativity preserving with $K(0_{+})\in(0,\infty).$ Let $T>0$ be fixed. For $N\in\mathbb{N}$, we define the corresponding approximations $\widehat{X}^{1},\widehat{X}^{2}$ and denote by $\xi^{1},\xi^{2}$ the auxiliary processes that arise in their construction. Note that we have by continuity of $K$ and $g$ we have
@@
Hence we can represent them in compact form as
@@
We show by induction on $k$ that $\widehat{X}^{1}_{t}\leq\widehat{X}^{2}_{t}$ on $(0,t_{k}]$. For $k=1,$ we have per assumption $\widehat{X}^{1}_{t}=\widetilde{g}_{1}(t)\leq\widetilde{g}_{2}(t)=\widehat{X}^{2}_{t}$ for $t\in(0,t_{1})$. Using the Comparison Principle for SDEs, see [27, Chapter VI, Theorem 1.1], we obtain $\widehat{X}^{1}_{t_{1}}=\xi_{t_{1}-}^{1,1}\leq\xi_{t_{1}-}^{2,1}=\widehat{X}_{t_{1}}^{2}$. Suppose that the result is true for some $k\geq 1$. Then, we have for all $j\in\{1,\dots,k\}$, by construction
@@
By induction, we get for $k=N$ that $\mathbb{P}[\widehat{X}^{1}_{T}\leq\widehat{X}^{2}_{T}]=1$. By (8) we conclude $\mathbb{P}[X^{1}_{T}\leq X^{2}_{T}]=1$. Since $T>0$ was arbitrary and $X$ has continuous sample paths on $(0,\infty)$, we obtain the desired comparison principle. Step 3: Let us now prove the assertion when $K(0_{+})=+\infty$ and $K$ is completely monotone. Using the Bernstein theorem, see [35, Theorem 4.8], we find a Borel measure $\mu$ on $\mathbb{R}_{+}$ such that
$K(t)=\int_{\mathbb{R}_{+}}\mathrm{e}^{-\rho t}\,\mu(\mathrm{d}\rho),\quad t\in(0,\infty).$
Define, for $H>0$, the approximation $K^{H}\colon\mathbb{R}_{+}\longrightarrow\mathbb{R}_{+}$ by
$K^{H}(t)\coloneqq\int_{[0,H]}\mathrm{e}^{-\rho t}\,\mu(\mathrm{d}\rho).$
Then $0\leq K^{H}(t)\leq K^{H}(0)$ and $|(K^{H})^{\prime}(t)|\leq HK^{H}(0)$. Hence, it satisfies Assumption 2.2 with $\gamma=1/2$ and therefore

### v2

@@ Abstract.
In this work, we establish a comparison principle for stochastic Volterra equations with respect to the initial condition and the drift $b$ applicable to a wide class of Volterra kernels and input curves $g$ that may be singular at zero. The latter appear, e.g., in the study of Markovian lifts for such Volterra equations. For completely monotone kernels, our result holds without any further restrictions, while for not completely monotone kernels, it is shown that such a principle fails unless the drift is additionally monotone. As a side-product of our results, we also complement the literature on the weak existence of continuous nonnegative solutions, which covers the rough Cox-Ingersoll-Ross process with singular initial conditions.
@@
Keywords: stochastic Volterra process; comparison principle; rough Cox-Ingersoll-Ross process; splitting method
@@ 1. Introduction
In many applications, one assumes that the Volterra kernel is given by the fractional Riemann-Liouville kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha\in(1/2,2)$. Such equations have gained increased attention in mathematical Finance for the modelling of rough volatility, see [8, 14, 15, 16, 19].
The weak existence of a solution was established in [4] whenever $b,\sigma$ are continuous with linear growth, $g$ is constant and $K$ satisfies an additional regularity condition given below. An extension to processes with jumps and more general $g$ is discussed in [1]. When $b$ and $\sigma$ are Lipschitz continuous, the existence of a unique strong solution was shown in [2, Theorem A.1], see also [9, 28]. For non-Lipschitz coefficients, the pathwise uniqueness of solutions is more subtle and yet not fully understood. For the fractional Riemann-Liouville kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha\in(1/2,1)$, the method proposed in [25] gives the existence and uniqueness of strong solutions whenever $\sigma$ is $\eta$-Hölder continuous and $\alpha\eta>\frac{1}{2}$. An extension of this method towards general drifts and non-convolution equations was recently discussed in [27]. Most recently, in [22], the author proves the weak existence and uniqueness in law for (1) provided that $b,\sigma$ are uniformly continuous, $\sigma$ is non-degenerate, and some technical condition on the modulus of continuity holds. In contrast to the existing literature, see also [26, 27], in this work we focus on continuous solutions where $g$ is not regular but belongs to $g\in L_{\mathrm{loc}}^{p}(\mathbb{R}_{+})\cap C((0,\infty))$ with an additional growth condition at the origin. Such functions $g$ that are singular in $t=0$ appear, e.g., from Markovian lifts where the initial condition does not belong to the domain of the corresponding projection operator, see e.g. [22, Theorem 3.3]. Thus, our Theorem 2.7 naturally complements the results obtained therein.
In the main part of this work, we study comparison principles for solutions of (1). More precisely, given initial data $(g_{i},K,b_{i},\sigma)$ with $i=1,2$, we provide sufficient conditions such that the corresponding solutions $X^{1},X^{2}$ satisfy
@@
While such comparison principles are well-known for classical ordinary differential equations (see [33]) and for Markovian stochastic equations (see [23] for the continuous case and [18] for the case of continuous-state branching processes), much less is known for their Volterra counterparts. For instance, comparison principles for deterministic Volterra equations have been studied in [21, Chapter 12.1], while the first results on stochastic Volterra equations were obtained in [17, 31] for general drifts and restrictive assumptions on the diffusion which exclude our convolution setting (1).
@@
Based on the splitting method, Alfonsi studied in [7] a numerical approximation of stochastic Volterra equations applicable to $g(t)=x$ constant, globally Lipschitz continuous coefficients $b,\sigma$, and $K$ being either non-negativity preserving with $K(0)>0$ or completely monotone. The latter work includes, to our knowledge, the first general comparison principle for equations of the form (1). This splitting method has also been recently applied in [6] to prove the existence and uniqueness of nonnegative solutions of stochastic Volterra equations with jumps for non-Lipschitz continuous coefficients and regular Volterra kernels $K\in C^{2}(\mathbb{R}_{+})$. The reader interested in other numerical schemes for stochastic Volterra processes may consult [3, 5] and the references therein. In this work, we first extend Alfonsi’s numerical approximation [7] to non-constant $g$ with a particular focus on possibly singular functions, and time-dependent coefficients. Based on this approximation, we obtain a comparison principle that covers such cases, including those that are not necessarily Lipschitz continuous. Afterwards, we focus on kernels $K$ that are not completely monotone, hence going beyond the framework discussed in [7].
The first key obstacle when dealing with stochastic Volterra equations stems from the absence of the semimartingale property. Indeed, while the classical methods described in [23, Chapter VI] are deeply based on the Itô formula, in our setting an analogous approach is only applicable when $K$ is sufficiently regular in the sense that $K\in W^{1,2}_{\mathrm{loc}}(\mathbb{R}_{+})$. The latter excludes, e.g., the Riemann-Liouville kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha\in(1/2,3/2)$. To address this issue, we extend the splitting method proposed in [7] to time-dependent coefficients and approximate the solution via solutions of classical stochastic differential equations. Observing that the approximation satisfies the comparison principle, we may then pass to the limit. This approach covers $K$ being either non-negativity preserving with $K(0)>0$ or completely monotone. In particular, it covers the fractional kernel with $\alpha\in(1/2,1)$. On the other side, if $K$ is neither non-negativity preserving nor completely monotone, a comparison principle is, in general, not valid as demonstrated in Example 3.1. The latter is, e.g., the case for the fractional kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha>1$. Such a failure of the comparison principle is not surprising since solutions of (1) are typically not Markov processes and hence it is not clear if the past trajectory entering the equation via $b,\sigma$ violates the ordering of solutions. To exclude such a possibility, we suppose that the drift is additionally monotone. Then, we first consider the case of regular kernels $K\in W_{\mathrm{loc}}^{1,2}(\mathbb{R}_{+})$ for which a semimartingale representation is available and a modification of the classical comparison principle for SDEs can be used to deduce the comparison principle. Afterwards, we extend these results to general $K\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})\cap C^{1}((0,\infty))$ with $K^{\prime}\geq 0$ or $K^{\prime}\leq 0$ by approximation.
This work is organised as follows. In Section 2 we treat the case of completely monotone kernels, where we first prove a comparison principle for Lipschitz continuous $b,\sigma$. Then we prove the existence of continuous weak solutions for continuous $b,\sigma$ with at most linear growth, and finally derive the corresponding comparison principle in this case. As an application, we conclude Section 2 with a result on the weak existence of continuous nonnegative solutions of (1). Afterwards, in section 3, we study similar results for the case of not completely monotone kernels. Some technical results on the existence, uniqueness, and sample path regularity of solutions to stochastic Volterra equations are collected in Appendix A, while the convergence of the splitting method is studied in Appendix B.
2. Non-negativity preserving case
In this section, we study the case where the Volterra kernel preserves nonnegativity. As a first step, let us introduce the minimal conditions on the input data $(g,K,b,\sigma)$ that will be used for the construction of continuous weak solutions.
Assumption 2.1.
There exist $C>0$, $\xi\in[0,1]$ such that
$|b(t,x)|+|\sigma(t,x)|\leq C(1+|x|)^{\xi}$
  1. (i)
  2. (ii)
@@ 2. Non-negativity preserving case
    $g\in L^{q}_{\mathrm{loc}}(\mathbb{R}_{+})$ for some $q\in(2,\infty)$, $\xi\in(0,1]$, there exists $\eta>0$ such that $K\in L^{2+\eta}_{\mathrm{loc}}(\mathbb{R}_{+})$ and there exist $\gamma\in(0,1/2]$ and for each $T>0$ a constant $c(T,\eta)$ such that for $h\in[0,1]$
    Moreover, the value $q$ satisfies
    (2) $q>2\xi\frac{1+\frac{1}{\eta}}{\gamma+\frac{1}{2}\frac{\eta}{2+\eta}}.$
The above assumption allows us to assume without loss of generality that for any solution $X$ of (1), XgX-g is continuous, c.f. Proposition A.2. Note that such a condition is satisfied by the fractional Riemann-Liouville kernel as demonstrated in the next example.
@@
Let $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha\in(1/2,1)$. Then $2+\eta<\frac{1}{1-\alpha}$, $\gamma=\alpha-1+\frac{1}{2+\eta}$. Hence (2) takes the particular form $q>2\xi\frac{1+1/\eta}{\alpha-1/2}$. By letting $\eta\nearrow\frac{2\alpha-1}{1-\alpha}$, we see that this condition is satisfied whenever
$q>\frac{\xi\alpha}{\left(\alpha-\frac{1}{2}\right)^{2}}.$
Next, to prove the desired comparison principles, let us introduce some concepts that allow us to compare the solutions. In this section, we focus on Volterra kernels that preserve nonnegativity in the sense of the following definition.
@@
    (3) $\sum_{\ell=1}^{k}x_{\ell}K(t_{k}-t_{\ell})\geq 0,\ \ \forall k\in\{1,\dots,N\}\Longrightarrow\sum_{\ell=1}^{N}\mathbbm{1}_{\{t_{\ell}\leq t\}}x_{\ell}K(t-t_{\ell})\geq 0,\quad\forall t\geq 0.$
Here we use the convention $\sum_{\varnothing}\coloneqq 0$. This condition was recently introduced and studied in [7], and allows us to show that the Volterra kernel preserves the order structure subject to the discretisation of the process via the splitting method. Finally, to obtain the comparison principle for two different input data $(g_{1},K,b_{1},\sigma)$ and $(g_{2},K,b_{2},\sigma)$, we use the following definition.
Definition 2.4.
Let $(g_{i},K,b_{i},\sigma)$ with $i=1,2$ satisfy Assumption 2.1, and suppose that $K\colon(0,\infty)\longrightarrow\mathbb{R}_{+}$ is continuous and preserves non-negativity. We say that $(g_{1},K,b_{1},\sigma)$ and $(g_{2},K,b_{2},\sigma)$ are comparable, if the following conditions hold:
@@
    There exist $h_{1},h_{2}\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})$ and $\widetilde{g_{1}},\widetilde{g}_{2}\in C((0,\infty))$ such that
    (4) $g_{i}(t)=\widetilde{g}_{i}(t)+\int_{0}^{t}K(t-s)h_{i}(s)\,\mathrm{d}s,\qquad i=1,2,$
    and $\widetilde{g}_{2}-\widetilde{g}_{1}\geq 0$ is non-decreasing on $(0,\infty)$.
@@
    $b_{1}(t,x)+h_{1}(t)\leq b_{2}(t,x)+h_{2}(t)$ holds for each $t\geq 0$ and $x\in\mathbb{R}$.
This class of comparable input data is sufficiently rich to cover the most frequently used examples from applications, as shown at the end of this section. The following is our first result for non-negativity preserving kernels.
@@
Suppose that $(g_{i},K,b_{i},\sigma)$ with $i=1,2$ are comparable in the sense of Definition 2.4. Assume that there exists $\delta>0$ with
(5) $\int_{0}^{T}\overline{g}_{i}(t)^{2+\delta}(1+h_{i}(t)^{2})\,\mathrm{d}t$ $<\infty,\qquad i=1,2,$
where $\overline{g}_{i}(t)=\sup_{s\in[t,T]}|g_{i}(s)|$, and that there exists $C>0$ such that for a.a. $t\geq 0$ and $x,y\in\mathbb{R}$,
@@
Let $X^{1},X^{2}$ be the unique strong solutions of (1) with input data $(g_{i},K,b_{i},\sigma)$, $i=1,2$. Then $\mathbb{P}[X^{1}_{t}\leq X^{2}_{t},\,t>0]=1$.
@@
First note that equation (1) has unique strong solutions $X^{1},X^{2}$ with sample paths in $L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$ due to Proposition A.1. Moreover, given Assumption 2.1, it follows from Proposition A.2 that $X^{1}-g_{1}$, $X^{2}-g_{2}$ have a modification with continuous sample paths. Hence $X^{1},X^{2}$ are continuous on $(0,\infty)$. Finally, define
@@
Hence $\widetilde{b}_{1}\leq\widetilde{b}_{2}$ and (6) still holds with the same $C$, while the linear growth constant becomes $|\widetilde{b}_{1}(t,x)|+|\widetilde{b}_{2}(t,x)|\leq\widetilde{C}(t)(1+|x|)$ with $\widetilde{C}(t)=C+|h_{1}(t)|+|h_{2}(t)|$, so that $\widetilde{C}\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$. The proof is divided into three steps.
Step 1: We introduce a similar splitting method to Alfonsi [7], but now applied for non-constant possibly singular at zero $g$ and time-dependent coefficients. Suppose first that $K$ is continuous with $K(0_{+})>0$. Let $T>0$, $N\in\mathbb{N}$ be fixed and $t_{k}=kT/N,$ $k\in\{0,1,\dots,N\},$ be an equidistant grid of $[0,T]$ with step size $T/N.$ Recursively, we define the process $(\widehat{X}^{i}_{t})_{t\in(0,T]}$ by setting
@@
Thus it suffices to prove that $\widehat{X}_{t}^{1}\leq\widehat{X}_{t}^{2}$ a.s. for $t\in(0,T]$.
Step 2: We consider the case where $K$ is continuous and non-negativity preserving with $K(0_{+})\in(0,\infty).$ Let $T>0$ be fixed. For $N\in\mathbb{N}$, we define the corresponding approximations $\widehat{X}^{1},\widehat{X}^{2}$ and denote by $\xi^{1},\xi^{2}$ the auxiliary processes that arise in their construction. Note that we have by continuity of $K$ and $g$
@@
Hence, we can represent them in compact form as
@@
We show by induction on $k$ that $\widehat{X}^{1}_{t}\leq\widehat{X}^{2}_{t}$ on $(0,t_{k}]$. For $k=1,$ we have per assumption $\widehat{X}^{1}_{t}=\widetilde{g}_{1}(t)\leq\widetilde{g}_{2}(t)=\widehat{X}^{2}_{t}$ for $t\in(0,t_{1})$. Using the Comparison Principle for SDEs, see [23, Chapter VI, Theorem 1.1], we obtain $\widehat{X}^{1}_{t_{1}}=\xi_{t_{1}-}^{1,1}\leq\xi_{t_{1}-}^{2,1}=\widehat{X}_{t_{1}}^{2}$. Suppose that the result is true for some $k\geq 1$. Then, we have for all $j\in\{1,\dots,k\}$, by construction
@@
By induction, we get for $k=N$ that $\mathbb{P}[\widehat{X}^{1}_{T}\leq\widehat{X}^{2}_{T}]=1$. By (7) we conclude $\mathbb{P}[X^{1}_{T}\leq X^{2}_{T}]=1$. Since $T>0$ was arbitrary and $X$ has continuous sample paths on $(0,\infty)$, we obtain the desired comparison principle.
Step 3: Let us now prove the assertion when $K(0_{+})=+\infty$ and $K$ is completely monotone. Using the Bernstein theorem, see [30, Theorem 4.8], we find a Borel measure $\mu$ on $\mathbb{R}_{+}$ such that $K(t)=\int_{\mathbb{R}_{+}}\mathrm{e}^{-\rho t}\,\mu(\mathrm{d}\rho)$ for $t\in(0,\infty)$. Define, for $H>0$, the approximation $K^{H}\colon\mathbb{R}_{+}\longrightarrow\mathbb{R}_{+}$ by $K^{H}(t)\coloneqq\int_{[0,H]}\mathrm{e}^{-\rho t}\,\mu(\mathrm{d}\rho)$. Then $0\leq K^{H}(t)\leq K^{H}(0)$ and $|(K^{H})^{\prime}(t)|\leq HK^{H}(0)$. Hence, it satisfies Assumption 2.1 with $\gamma=1/(2+\eta)$ and therefore